# Batch Acquisition Function Evaluations and Decouple Optimizer Updates for Faster Bayesian Optimization

### v1

@@ Abstract
Bayesian optimization (BO) efficiently finds high-performing parameters by maximizing an acquisition function, which models the promise of parameters. A major computational bottleneck arises in acquisition function optimization, where multi-start optimization (MSO) with quasi-Newton (QN) methods is required due to the non-convexity of the acquisition function. BoTorch, a widely used BO library, currently optimizes the summed acquisition function over multiple points, leading to the speedup of MSO owing to PyTorch batching. Nevertheless, this paper empirically demonstrates the suboptimality of this approach in terms of off-diagonal approximation errors in the inverse Hessian of a QN method, slowing down its convergence. To address this problem, we propose to decouple QN updates using a coroutine while batching the acquisition function calls. Our approach not only yields the theoretically identical convergence to the sequential MSO but also drastically reduces the wall-clock time compared to the previous approaches.
Codehttps://github.com/Kaichi-Irie/aaai2026-anon-exp/tree/submission
@@ 5 Benchmarking Experiments
Table 1 shows the benchmarking results. Based on the results, all methods achieved comparable (median) final objective values under the shared iteration cap. While C-BE occasionally outperformed Seq. Opt. for $D\in\{10,20\}$, we attribute this to seed variability and the median-of-minima aggregation rather than a systematic benefit of coupling, in contrast to the sizable iteration increase observed under C-BE. Meanwhile, as can be seen in the Iters. column, the L-BFGS-B iteration count drastically increased in C-BE. For example, C-BE required substantially more L-BFGS-B iterations than Seq. Opt. as $D$ grew, e.g., $68.8/21.5\!\approx\!3.2$ for $D=20$ and $87.0/26.5\!\approx\!3.3$ for D=40D{=}40. This observation is consistent with the discussion regarding the off-diagonal artifacts. These extra iterations can offset batching gains, making C-BE’s wall-clock time comparable to—and on some objectives worse than—Seq. Opt. By contrast, D-BE matches the iteration counts of the Seq. Opt. baseline while exploiting batched gradient evaluations, having reduced the runtime of the acquisition function optimization ($\sim\!1.5\times$ faster than Seq. Opt.). The same pattern holds on Sphere, Attractive Sector and Step Ellipsoidal from the COCO (BBOB) suite (hansen2009real; hansen2021coco); See Appendix C (Table 2) for more details.
@@ 6 Conclusion
Appendix A Pseudocode: Sequential Optimization (Seq. Opt.)
@@
Algorithm 2 presents the Seq. Opt. algorithm, which serves as a baseline: each restart maintains its own optimizer state, and no batching is performed across restarts. As a result, no cross-restart information is shared, and the per-restart curvature is preserved by construction.
Appendix B BFGS Convergence on Separable Sums
We verify that the observed degradation under C-BE is not a limited-memory artifact: the same pattern appears with BFGS on Rosenbrock.
Figure 3: Same as Figure 1 but BFGS, i.e., memory is not limited, with B=3B{=}3, D=5D{=}5. (a) ground truth (block diagonal); (b) Seq. Opt.: structure preserved; (c) C-BE: dense off-diagonal couplings.
Figure 4: Same as Figure 1 but BFGS with B=10B{=}10, D=5D{=}5. Off-diagonal seems to be more serious with larger $B$.
Figure 5: Same as Figure 2 but BFGS. C-BE (B>1B{>}1) requires substantially more iterations than the Seq. Opt. baseline (B=1B{=}1 per restart).
We replicate the setup from Section 3 using BFGS with $B\in\{1,2,5,10\}$ restarts on D=5D{=}5 Rosenbrock function. Figures 34 show that C-BE again injects off-diagonal mass into the inverse Hessian approximation that should be zero for a separable sum objective, while Seq. Opt. preserves the block-diagonal structure. Consistently, iteration counts increase markedly under C-BE relative to Seq. Opt. (Figure 5).
These observations indicate that the failure arises from applying QN methods updates to a coupled yet separable problem (shared state across restarts), rather than from the limited-memory approximation.
Appendix C Full Benchmark Results
Objective ($D$) Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
Sphere (5) Seq. Opt. $4.026\times 10^{-5}$ $196.9$ $\mathbf{7.0}$
@@ Appendix C Full Benchmark Results
Table 2: Benchmarking results of BO ($300$ trials) using L-BFGS-B ($m=10$) with $B=10$ restarts on four objectives: Sphere, Attractive Sector (AS), Step Ellipsoidal (SE), and Rastrigin, each at $D\in\{5,10,20,40\}$. The Objective ($D$) column lists the objective name and dimensionality (e.g., Sphere (10) means $D=10$). The termination criterion of L-BFGS-B is $200$ iterations or $\|\nabla\alpha(\mathbf{x})\|_{\infty}\leq 10^{-2}$. Each entry is the median over 2020 independent runs with different seeds. Best Value: $\min$ objective among the $300$ trials minus the best objective over all runs. Runtime (s): total wall-clock time for BO. median L-BFGS-B iteration count over $300$ trials $\times$ $10$ restarts. Lower is better.
This appendix complements the main text (which reported only Rastrigin) by providing the complete results for Sphere, Attractive Sector, Step Ellipsoidal, and Rastrigin at $D\in\{5,10,20,40\}$ in Table 2. Experimental settings are identical to those stated in the caption and the main text.
Across objectives and dimensions, C-BE degrades notably in higher dimensions both in solution quality and in optimizer iterations. At D=40D{=}40, C-BE yields larger Best Value than D-BE by $25.4\%$ on Attractive Sector, $16.5\%$ on Step Ellipsoidal, and $31.3\%$ on Rastrigin (Sphere is roughly on par: $2.8\%$). [^2] Percentages computed as $(\textsc{C-BE}-\textsc{D-BE})/\textsc{D-BE}$ using Table 2. Its iteration counts also inflate sharply: at D=40D{=}40, C-BE requires about $3.3{\times}\!\sim\!29{\times}$ more L-BFGS-B iterations than D-BE ($3.3{\times}\!\sim\!34{\times}$ vs. Seq. Opt.), with the largest gap on Attractive Sector (102.2 vs. 3.5 and 3.0).

### v2

@@ Abstract
Bayesian optimization (BO) efficiently finds high-performing parameters by maximizing an acquisition function, which models the promise of parameters. A major computational bottleneck arises in acquisition function optimization, where multi-start optimization (MSO) with quasi-Newton (QN) methods is required due to the non-convexity of the acquisition function. BoTorch, a widely used BO library, currently optimizes the summed acquisition function over multiple points, leading to the speedup of MSO owing to PyTorch batching. Nevertheless, this paper empirically demonstrates the suboptimality of this approach in terms of off-diagonal approximation errors in the inverse Hessian of a QN method, slowing down its convergence. To address this problem, we propose to decouple QN updates using a coroutine while batching the acquisition function calls. Our approach not only yields the theoretically identical convergence to the sequential MSO but also drastically reduces the wall-clock time compared to the previous approaches. Our approach is available in GPSampler in Optuna, effectively reducing its computational overhead.
Codehttps://github.com/Kaichi-Irie/faster-batched-acqf-opt-experiments/tree/submission
@@ 1 Introduction
Notice that our approach has been successfully merged into GPSampler in Optuna, significantly accelerating its runtime.
@@ 5 Benchmarking Experiments
Table 1 shows the benchmarking results. Based on the results, all methods achieved comparable (median) final objective values under the shared iteration cap. As can be seen in the Iters. column, the L-BFGS-B iteration count drastically increased in C-BE. For example, C-BE required substantially more L-BFGS-B iterations than Seq. Opt. as $D$ grew, e.g., $68.8/21.5\!\approx\!3.2$ for $D=20$ and $87.0/26.5\!\approx\!3.3$ for D=40D{=}40. This observation is consistent with the discussion regarding the off-diagonal artifacts. These extra iterations can offset batching gains, making C-BE’s wall-clock time comparable to—and on some objectives worse than—Seq. Opt. By contrast, D-BE matches the iteration counts of the Seq. Opt. baseline while exploiting batched gradient evaluations, having reduced the runtime of the acquisition function optimization ($\sim\!1.5\times$ faster than Seq. Opt.). The similar pattern holds on Sphere, Attractive Sector and Step Ellipsoidal from the COCO (BBOB) suite (hansen2009real; hansen2021coco); See Appendix C (Table 2) for more details. Note that while C-BE occasionally outperformed Seq. Opt. for $D\in\{10,20\}$, we attribute this to the difference in the convergence behavior rather than a systematic benefit of coupling. More specifically, the acquisition function optimization for C-BE is mostly terminated by the gradient tolerance but not the iteration count, meaning that the optimization results for C-BE are not premature. This leads to very different optimization results between C-BE and Seq. Opt., making it possible for C-BE to sometimes outperform Seq. Opt.
@@ 6 Conclusion
Appendix A Sequential Optimization (Seq. Opt.)
@@
Algorithm 2 presents the Seq. Opt. algorithm, which serves as a baseline. Each restart maintains its own optimizer state, and no batching is performed across restarts. As a result, no cross-restart information is shared, and the per-restart curvature is preserved by construction.
Appendix B Off-Diagonal Artifact Effect on BFGS
Figure 3: Contour maps of the inverse Hessian (Left) and its approximations by Seq. Opt. (Center) and C-BE (Right). This figure follows the same setup as Figure 1 except that BFGS, i.e., the memory is not limited, is used instead of L-BFGS-B.
Figure 4: Contour maps of the inverse Hessian (Left) and its approximations by Seq. Opt. (Center) and C-BE (Right). This figure follows the same setup as Figure 1 except that BFGS, i.e., the memory is not limited, is used instead of L-BFGS-B, and $B=10$ is used instead of $B=3$. Off-diagonal artifacts are more prominent for a large $B$.
Figure 5: Convergence speed of C-BE when using BFGS. This figure follows the same setup as Figure 2 except that BFGS is used instead of L-BFGS-B. As the number of restarts $B$ increases, the convergence of C-BE requires substantially more iterations.
This section verifies that the observed degradation under C-BE does not stem from limiting the memory of BFGS. The setup in this section follows Section 3. Figures 3 and 4 show that C-BE injects off-diagonal mass into the inverse Hessian approximation, which should be zero, while Seq. Opt. preserves the block-diagonal structure. Consistently, iteration counts increase markedly under C-BE relative to Seq. Opt. as seen in Figure 5. These observations confirm that the off-diagonal artifacts are caused by coupling QN updates for a separable problem rather than by limiting the memory.
Appendix C Full Benchmarking Results
Objective Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
5DD Sphere Seq. Opt. $4.026\times 10^{-5}$ $196.9$ $\mathbf{7.0}$
@@ Appendix C Full Benchmarking Results
Table 2: The benchmarking results of BO with $300$ trials using L-BFGS-B with the memory size of $m=10$ and $B=10$ restarts on the $D$-dimensional functions ($D=5,10,20,40$). We picked four objectives: Sphere, Attractive Sector (AS), Step Ellipsoidal (SE), and Rastrigin. The termination criterion of L-BFGS-B is $200$ iterations or $\|\nabla\alpha(\mathbf{x})\|_{\infty}\leq 10^{-2}$. Each column shows the median over 2020 independent runs, each with a different random seed. Best Value: the minimum objective value among the 300 trials minus the best objective value over all runs. Runtime (s): the total wall-clock time for BO. Iters.: the median of the L-BFGS-B iteration count over $300$ Trials $\times 10$ restarts. Lower is better.
Table 2 provides the complete results for Sphere, Attractive Sector, Step Ellipsoidal, and Rastrigin. Note that the setup follows Section 5. The results show that C-BE degrades notably in higher dimensions both in solution quality and in optimizer iterations. For example, C-BE yields larger Best Value in 40DD than D-BE by $25.4\%$ on Attractive Sector, $16.5\%$ on Step Ellipsoidal, $31.3\%$ on Rastrigin, and $2.8\%$ on Sphere  [^2] Percentages computed as $(\textsc{C-BE}-\textsc{D-BE})/\textsc{D-BE}$ using Table 2. . The iteration counts of C-BE also inflate sharply at D=40D{=}40, requiring about $3.3{\times}\!\sim\!29{\times}$ more L-BFGS-B iterations than D-BE. The largest gap is observed on Attractive Sector (102.2 vs. 3.5).